OpenStudy (anonymous):

A plant 3 centimeters tall is placed by a window. If it grows 4 centimeters a day, how many days will it take to be 51 centimeters tall? A.12 B.17 C.12.75 D.13

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

51 = 3 + 4t solve for t

OpenStudy (anonymous):

-3 on both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, 4t=47?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 4t = 48

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Oh.

OpenStudy (whpalmer4):

You could think of this as dividing the amount the plant needs to grow (51 cm - 3 cm=48 cm) into however many days of growth at 4 cm/day. What is 48 / 4?

OpenStudy (anonymous):

So...t=12?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

Yaya!

OpenStudy (anonymous):

lool

OpenStudy (whpalmer4):

now check your work: 3 cm+ 12 days * 4 cm/day = 3 cm + 48 cm = 51 cm

OpenStudy (anonymous):

Yaya

OpenStudy (anonymous):

who says that stuff nowadays...

OpenStudy (anonymous):

What? Yaya?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Well then..

OpenStudy (anonymous):

Try this...

OpenStudy (whpalmer4):

the charming @hadleydarcym does, that's good enough for me! ;-)

OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A.w + 5 = 116 B.4w + 10 = 116 C.2w + 5 = 116 D.2w + 10 = 116

OpenStudy (anonymous):

Awe! Thank you!

OpenStudy (whpalmer4):

now, who confuses "awe" and "aw"? unfortunately, same answer :-)

OpenStudy (anonymous):

Hmm..Hoekay. (:

OpenStudy (whpalmer4):

what's the formula for the perimeter of a rectangle?

OpenStudy (anonymous):

-_-

OpenStudy (whpalmer4):

what is the perimeter?

OpenStudy (whpalmer4):

it's just the length of all the sides, right?

OpenStudy (anonymous):

Peremeter is adding up the lengths of all sides.

OpenStudy (anonymous):

...yes

OpenStudy (whpalmer4):

good. so, we have a rectangle, and it has dimensions l for length and w for width. can you write the formula for perimeter of that rectangle? it's just all the sides added up...

OpenStudy (anonymous):

here you know that 2w + 2l = 116 and that l = (w+5)

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

sub for l you get 2(w+5) +2w = 116

OpenStudy (anonymous):

...2w+10+2w=116

OpenStudy (anonymous):

than expand and simplify 2w + 10 +2w = 116 =) you can do the rest now..common @hadleydarcym

OpenStudy (anonymous):

Combine like terms?

OpenStudy (whpalmer4):

\[P = l + l+ w+w\] and as @mebs pointed out, \(l = w+5\) because the length is 5 feet more than the width. \[P = (w+5) + (w+5) + w + w\]and that whole thing we know to be equal to 116 so \[116 = (w+5)+(w+5)+w+w\]

OpenStudy (whpalmer4):

combining like terms is good.

OpenStudy (whpalmer4):

remember, you're just trying to get it to look like one of your answer choices, so take a look at the list after each manipulation!

OpenStudy (anonymous):

Its d. Right?

OpenStudy (whpalmer4):

p.s. as long as I'm correcting people's word choices here, it's "come on", not "common" :-) @hadleydarcym sorry, that's incorrect. don't you have 4 w's in your equation?

OpenStudy (anonymous):

It auto submitted it. I was timed...

OpenStudy (anonymous):

Thanks though.

OpenStudy (whpalmer4):

\[116 = (w+5) + (w+5) + w + w\]\[116 = w + 5 + w + 5 + w + w\]\[116 = 4*w + 10\]

OpenStudy (whpalmer4):

I'm more interested in you learning how to do it, not so concerned about getting any one question done in time...

OpenStudy (whpalmer4):

can you solve that equation for w?

OpenStudy (anonymous):

4w+10. Hmm.

OpenStudy (anonymous):

YAYA!

OpenStudy (whpalmer4):

\[116 = 4w+10\]You want to get the w term all by itself on one side...

OpenStudy (anonymous):

Yay

OpenStudy (whpalmer4):

how about subtracting 10 from both sides, what does that give you?

OpenStudy (anonymous):

4w=106?

OpenStudy (anonymous):

And then divide them both by 4?

OpenStudy (anonymous):

Yaya

OpenStudy (anonymous):

Alright now @mebs . I see you like my words...

OpenStudy (anonymous):

=D

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

lool cya yw

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

What the frack is that supposed to mean?

OpenStudy (anonymous):

what do you mean what whats supposed to mean? =?

OpenStudy (anonymous):

The whatever you said..like...lool cyl yw? What the frack?

OpenStudy (anonymous):

its translated into your language as: haha your welcome

OpenStudy (whpalmer4):

what, did you lose your magic decoder ring?

OpenStudy (anonymous):

yea what is the meaning of this!!?

OpenStudy (anonymous):

I don't even know. -_-

OpenStudy (anonymous):

well then -.-'''

OpenStudy (anonymous):

And whaaaatttt?

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

that supposed to be seat pouring down your forehead after doing tedious math problems.. that rarely happens to me...

OpenStudy (anonymous):

sweat**

OpenStudy (whpalmer4):

the silliness factor here is unbounded, I think :-)

OpenStudy (anonymous):

Hahaha. I'm just sitting here...Listening to all of this. (:

OpenStudy (anonymous):

Well...Reading

OpenStudy (anonymous):

@whpalmer4 yea math people are hilarious haha

OpenStudy (whpalmer4):

often without even trying to be :-) or as a friend used to say "funny ha ha or funny strange" :-)

OpenStudy (anonymous):

you should see my math prof when he's trying to crack a joke ...

OpenStudy (anonymous):

Haha

OpenStudy (whpalmer4):

my old calculus teacher, bless his heart, always mispronounced asymptote...he said as_mytote (_ represents a letter which will cause the openstudy censor to change the word)

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I get really scared because he might be giving us some really hard problems!! It's strange, I like how he's Russian and tries to speak English its too funny.

OpenStudy (anonymous):

Wow. Just. Wooowwwww. -_-

OpenStudy (whpalmer4):

as long as we're talking math silliness, I think you should both go listen to this very fine song brought to you by a former math professor named Tom Lehrer: http://www.youtube.com/watch?v=qU_j5cQ2sfQ

OpenStudy (anonymous):

When he can't find the right word in English to explain himself, he swears in Russian!!!! It's absolutely hilarious...not only that but there are only like 6 people in his class including me haha..

OpenStudy (anonymous):

is math humor to cheesy for you =S @hadleydarcym haha

OpenStudy (anonymous):

Nahh. Lol. It's fine. (:

OpenStudy (anonymous):

I am joking @hadleydarcym ... I really am. Now get back to work....

OpenStudy (anonymous):

Work? What is this work concept you speak of? XD

OpenStudy (anonymous):

there...your getting the hang of it...Get to work!!!

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

i am confuseddddedddeeddd